Mixing a 10 percent solution with a 60 percent solution to get 100 grams of a 50 percent solution. How much of each do i need?
Answer by ewatrrr(24785)


Hi,
to get 100 grams of a 50 percent solution...
Let x represent the amount of the 60% solution, (assigning x to the stronger solution)
then (100g - x)would represent the amount of the 10% solution
Mixing the 10 percent solution with the 60 percent solution****
.10(100g - x) + .60x = .50*100g
solving for x
.10*100g - .10x +.60x = .50*100g
.50x = .40*100g
x = 80g, the amount of the 60% solution, 20g of the 10% solution (100g-80g)
checking our amswer
.10*20g + .60*80g = 2g + 48g = 50g = .50*100g